For a given week, a random sample of 30 hourly employees selected from a very large number of employees in a manufacturing firm has a sample mean wage of x̅ = $180.00, with a sample standard deviation of
1
Expert's answer
2020-12-28T13:39:06-0500

The provided sample mean is "\bar{X}=180" and the population standard deviation is "\sigma=14." The required confidence level is "95\%," and the significance level is "\alpha=0.05." Based on the provided information, the critical "z"-value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The 95% confidence for the population mean "\mu" is computed using the following expression


"CI=(\bar{X}-z_c\times\dfrac{\sigma}{\sqrt{n}},\bar{X}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(180-1.96\times\dfrac{14}{\sqrt{30}},180+1.96\times\dfrac{14}{\sqrt{30}})"

"=(174.99, 185.01)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "174.99<\mu <185.01," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(174.99, 185.01)."